OpenStudy (anonymous):

john has two rectangular mirror that are similar. the lenght of the smaller miirror is 4 inches and width is 6 inches. the width of the bigger mirror is 2 feet. what is the lenght of the bigger mirror?

OpenStudy (anonymous):

(4 inches/ x inches)=(6 inches/2*12 inches) 4/x=6/24 cross multiply and you get 4*24=6x 96=6x x=16 the length is 16 inches or a foot and four inches